# Anant-Net: Breaking the Curse of Dimensionality with Scalable and Interpretable Neural Surrogate for High-Dimensional PDEs

### v2

@@ 2.1 Anant-Net’s Data Structures: Balancing Training Efficiency and Testing Flexibility
For the Anant-Net architecture, it is important to highlight the non-trivial differences between the training and testing procedures. During training, the data structure follows a separable (or tensor-array) format, as described in Algorithm 1. In this format, only $B$ selected dimensions are active, meaning they exhibit active gradients, while the remaining $(d-B)$ dimensions remain fixed in an arbitrary $d$-dimensional problem. This tensor-based format offers enhanced memory efficiency, which helps prevent memory bottlenecks even in problems with extremely large dimensionality. Furthermore, the tensor-based training framework enables the data to be handled as hypercubes, fully leveraging the Anant-Net architecture described in Algorithm 2. However, using a tensor-array structure during the testing phase would render the model overly rigid, thereby limiting its applicability to more practical scenarios involving randomly selected dimensions. To address this, the testing phase employs a distinct approach, wherein the data structure adopts a vector-array format, consisting of randomly sampled testing points (and not a grid). This format allows unrestricted selection of dimensions from an arbitrary $d$-dimensional space during inference.
@@ 2.2 Theoretical Analysis of Computational Complexity
Here, we consider space complexity as the function of the number of trainable parameters in a given model architecture. For a fully-connected neural network of depth, $D$ (layers) and width, $W$ (neurons per layer), the space complexity of the model can be approximated as $\mathcal{O}(W^{2}D)$ assuming no bias terms. However, for a separable neural network (such as SPINN[28]) with $B$ body networks, the space complexity is given by $\mathcal{O}(B\tilde{W}^{2}\tilde{D})$ where $\tilde{D}<D\text{ and }\tilde{W}<W$. To ensure a fair comparison, we assume $\mathcal{O}(B\tilde{W}^{2}\tilde{D})\approx\mathcal{O}(W^{2}D)$ thereby keeping the representation capacity approximately the same across both architectures. However, for such separable architecture, the space complexity becomes exceedingly large as the complexity scales linearly with the dimension, $d$ of the PDE since $B=d$ leading to inefficiencies for high-dimensional problems. In comparison, Anant-Net also follows the separable architecture paradigm, with space complexity $\mathcal{O}(B\tilde{W}^{2}\tilde{D})$, but importantly with $B<<d$. As a result, the complexity of Anant-Net does not scale with dimensionality of the PDE, making it significantly more scalable.
@@ 2.3 Universal Approximation Theorem: Through the lens of Anant-Net
In this section we define the universal approximation capabilities of Anant-Net architecture solving high-dimensional PDEs. For clarity, the Anant-Net ansatz is formulated using Einstein summation convention rather than explicit tensor product notation ($\otimes$).
@@
(Anant-Net) Let X, Y be compact subsets of $\mathbb{R}^{d}$. With $u\in L^{2}(X\times Y).$ Then, for arbitrary $\varepsilon>0$ we can find a sufficiently large $r>0$ (which is the size of the embedding layer for all body networks) and neural networks $f_{j}$ and $g_{j}$ such that,
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{j=1}^{r}f_{j}g_{j}\biggr{\|}_{L^{2}(X\times Y)}<\varepsilon$ (4)
where, ^={^$\hat{\mathcal{B}}=\{\hat{\mathcal{B}}_{1},\hat{\mathcal{B}}_{2},\dots\hat{\mathcal{B}}_{k}\}$ is the set of sets consisting of active dimensions where at any given step-$k$, ^$\hat{\mathcal{B}}_{k}=\{\hat{\mathcal{B}}_{k}^{1},\hat{\mathcal{B}}_{k}^{2}\}=\{x_{p},x_{q}\}$ is the set of active dimensions, $\{X$\^$\hat{\mathcal{B}}_{k}^{1}\}$ and $\{Y$\^$\hat{\mathcal{B}}_{k}^{2}\}$ is the set of inactive dimensions. Here, $X=\{x_{1},x_{2},\dots x_{p}\dots x_{d/2}\}$, $Y=\{x_{d/2+1},\dots x_{q}\dots x_{d}\}$.
@@ 2.4 Effect of Preconditioning on Training Dynamics
where $f^{(\theta_{k})}(X_{k})=\sum_{i=1}^{n_{k}}w_{2,k}^{i}(w_{1,k}X_{k}+b_{1})=\sum_{i=1}^{n_{k}}w_{2,k}^{i}\phi_{i}(X_{k}),\leavevmode\nobreak\ k=1,2,3.$ The parameter $w^{i}_{j,k}$ denotes the weight associated with the $i^{\text{th}}$ neuron in the $j^{\text{th}}$ layer of the $k^{\text{th}}$ body network. In the above formulation, the basis functions are implicitly represented by the hidden layers of the body networks within the Anant-Net architecture, while the corresponding coefficients are extracted from the final output layer. This representation is analogous to the framework proposed by Cyr et al. [32], wherein a generic deep neural network is interpreted as a linear combination of learned basis functions. We consider the continuous formulation of the loss function, which serves as the analog of the discrete loss representation given in Equation (3), corresponding to the same PDE defined in Equation (2).
[…]
@@
where $H_{r}$ is the hessian computed at intermediate parameters between 00 and $r$. Here, for brevity a new notation $\phi_{i}(x)=\partial_{\theta_{i}}u(x;\theta_{0})$ is introduced with $\mathcal{L}\phi_{i}\in L^{2}(\Omega)$. The above formulation is substituted in Equation 6 that leads to the following simplification:
@@
[…]
@@
In the case where the matrix $\mathbb{A}$ is invertible, let $\kappa(\mathbb{A})$ denote its condition number. To achieve a specified error tolerance $\epsilon$, the number of gradient descent iterations required satisfies the relation $N(\epsilon)\propto 1/\ln(1-c/\kappa(\mathbb{A}))$ for some constant $10<c<$ 1. Consequently, as $\mathbb{A}$ becomes increasingly ill-conditioned, i.e., $\kappa(\mathbb{A})\to\infty$, the number of iterations $N(\epsilon)$ diverges, indicating a deterioration in convergence rate. To address this challenge in the training of physics-informed machine learning models, we employ an explicit preconditioning strategy for the gradients, as detailed below.
@@
$\widehat{\theta}_{r+1}=\widehat{\theta}_{r}-\eta H^{-1}\nabla L(\widehat{\theta}_{r})$ (9)
[…]
@@ 2.5 Anant-Net with Interpretable KAN
$\Phi_{l}=\begin{bmatrix}\phi_{l,1,1}(\cdot)&\phi_{l,1,2}(\cdot)&\dots&\phi_{l,1,n_{l}}(\cdot)\\ \phi_{l,2,1}(\cdot)&\phi_{l,2,2}(\cdot)&\dots&\phi_{l,2,n_{l}}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\phi_{l,n_{l+1},2}(\cdot)&\dots&\phi_{l,n_{l+1},n_{l}}(\cdot)\\ \end{bmatrix}$
@@
Figure 4: Schematic of 3-layered Kolmogorov Arnold Networks (KAN) with input, $\mathbf{x}_{0}=\{x_{1},x_{2}\}$ and output, $f(\mathbf{x})=(\Phi_{2}\circ\Phi_{1}\circ\Phi_{0})(\mathbf{x}_{0})$. Replacing the ANN (body networks) in the Anant-Net shown in Figure 2 by plugging KAN (body networks) results in Anant-KAN architecture.
@@ 3 Results
This section presents the empirical investigations performed using Anant-Net to solve high-dimensional linear and nonlinear PDEs over hypercubes with domain length $L>1$, where the volume increases with the number of dimensions. Specifically, we focus on the Poisson equation, sine-Gordon equation, and Allen-Cahn equation. A key challenge in such settings is the enforcement of boundary conditions, where boundary sampling significantly contributes to the complexity of the problem as the dimensionality increases. To evaluate the performance of the proposed architecture, we use the percentage relative $L_{2}$ error, defined as:
@@
Uncertainty quantification is performed by reporting the test error over 10 independent random realizations, with data points sampled from the high-dimensional domain. Unless otherwise specified, all experiments are executed on a T4 GPU with 15 GB RAM. To this end, we systematically compare the performance of Anant-Net with other state-of-the-art methods that also employ boundary sampling strategies, rather than enforcing boundary conditions directly. The weights for the individual loss terms is fixed for all the cases: $\lambda_{r}=1$ and $\lambda_{b}=15$ in Equation 3.
In addition to the architectural hyperparameters in deep neural networks, several algorithmic hyperparameters must be tuned to achieve optimal performance with the Anant-Net architecture. These include: the number of body networks ($B$), grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#$\#CG$), number of boundary/data grids (#$\#BG$), and the sampling frequency ($S_{\text{FREQ}}$). The number of body networks ($B$) directly influences the architecture of Anant-Net. For example, when $B=3$, the architecture consists of three body networks that process 3D cubes sampled from the high-dimensional space. Similarly, when $B=4$, four body networks are used to process 4D tesseracts from the high-dimensional domain. The sampling frequency ($S_{\text{FREQ}}$) directly affects the training process by determining how frequently a new set of dimensions is sampled during training. If $S_{\text{FREQ}}$ is too low, the model spends less time interacting with the newly sampled subspace, leading to poor generalization. Conversely, if $S_{\text{FREQ}}$ is too high, the model may overfit due to prolonged exposure to the same subspace. Finally, the grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#$\#CG$), and number of boundary/data grids (#$\#BG$) determine the size and structure of the dataset handled by Anant-Net. We will employed $B=3$ for all computational experiments unless stated otherwise.
Figure 5: Point-wise absolute testing error on randomly selected 3D cubes from 21D, 99D, and 300D cases for the (left) Poisson equation, (center) sine-Gordon equation, and (right) Allen–Cahn equation. Here, ”$\mathsf{ActiveDims}$: $\{p,q,r\}$” denotes the active dimensions in the $d$-dimensional space where gradients are non-zero, and $\mathrm{D}$\$\{p,q,r\}$ represents the inactive dimensions. $\mathrm{D}=\{1,2,\dots,d\}$ denotes the set of all dimensions.
@@ 3.1 High-dimensional Poisson’s equation
Dimension Architecture Optimizer ##\##Collocation ##\##Data
@@
In this section, we solve the high-dimensional Poisson’s equation in 21D, 99D, and 300D. For all cases, we use the Adam optimizer to train Anant-Net, followed by L-BFGS to further fine-tune the learning process. The benefits of preconditioning the gradients were discussed in previous sections, where we empirically validated the theory using (quasi-)Newton optimization with a linear version of Anant-Net. However, in this section, Anant-Net incorporates a non-linear activation function ($\tanh$) to solve the high-dimensional Poisson’s equation. Table 2 provides further details on the specific hyperparameters used. As suggested by the curse of dimensionality, the number of collocation and data samples required to solve high-dimensional PDEs on a hypercube grows exponentially with dimension. This poses a major challenge for existing methods, including state-of-the-art PINNs such as NTK-based PINNs [29], self-adaptive PINNs [39], etc., which struggle with scalability due to the computational burden of large sample volumes. Anant-Net addresses this issue efficiently. Table 2 summarizes the total number of collocation and boundary points used for solving Poisson’s equation in 21D, 99D, and 300D. The sample sizes are controlled via hyperparameters: $N_{C}$, #$\#CG$ for collocation points, and $N_{B}$, #$\#BG$ for boundary points. Figure 6 (left) indicates that boundary data sampling is the primary contributor to training time across all dimensions, highlighting a need to optimize this component. Meanwhile, Figure 6 (right) shows a mean relative $L_{2}$ testing error along with the standard deviation, evaluated on randomly selected test points from the high-dimensional space. Figure 5 (left) shows the absolute point-wise error on randomly selected 3D cubes from the 21D, 99D, and 300D cases. More details on the hyperparameters used for training is explained in APPENDIX B.

### v3

@@ 2.1 Anant-Net’s Data Structures: Balancing Training Efficiency and Testing Flexibility
For the Anant-Net architecture, it is important to highlight the non-trivial differences between the training and testing procedures. During training, the data structure follows a tensor-array format, as described in Algorithm 1. In this format, only $B$ selected dimensions are active, meaning they exhibit active gradients, while the remaining (dB)(d-B) dimensions remain fixed in an arbitrary $d$-dimensional problem. This tensor-based format offers enhanced memory efficiency, which helps prevent memory bottlenecks even in problems with extremely large dimensionality. Furthermore, the tensor-based training framework enables the data to be handled as hypercubes, fully leveraging the Anant-Net architecture described in Algorithm 2. However, using a tensor-array structure during the testing phase would render the model overly rigid, thereby limiting its applicability to more practical scenarios involving randomly selected dimensions. To address this, the testing phase employs a distinct approach, wherein the data structure adopts a vector-array format, consisting of randomly sampled testing points (and not a grid). This format allows unrestricted selection of dimensions from an arbitrary $d$-dimensional space during inference.
@@ 2.2 Theoretical Analysis of Computational Complexity
[…]
@@ 2.3 Universal Approximation Theorem: Through the Lens of Anant-Net
The Universal Approximation Theorem (UAT) provides a foundational guarantee for neural networks, stating that a feedforward neural network with a single hidden layer containing a sufficient number of neurons and a non-linear activation function can approximate any continuous function on a compact domain to arbitrary accuracy. This result highlights the theoretical expressivity of neural architectures and their applicability to a wide range of function approximation tasks, including the solution of PDEs.
In the context of scientific machine learning, UAT plays an important role in justifying the use of neural networks as surrogates or solution ansätze for PDEs. In this section, we analyze the universal approximation capabilities of the proposed Anant-Net architecture, specifically in the context of high-dimensional PDEs. Our goal is to illustrate, both theoretically and empirically, that Anant-Net retains sufficient expressivity to approximate complex solution manifolds that arise in such problems. To facilitate compact and efficient notation in our analysis, we adopt the Einstein summation convention in formulating the Anant-Net ansatz, rather than relying on explicit tensor product notation ($\otimes$).
@@
(Anant-Net) Let X, Y be compact subsets of $\mathbb{R}^{d}$. With $u\in L^{2}(X\times Y).$ Then, for arbitrary $\varepsilon>0$ we can find a sufficiently large $r>0$ (which is the size of the embedding layer for all body networks) and neural networks $f_{j}\triangleq f_{1,j}^{(\theta_{1})}$ and $g_{j}\triangleq f_{2,j}^{(\theta_{2})}$ such that,
$\frac{1}{|\hat{\mathcal{B}}|}\sum_{k=1}^{|\hat{\mathcal{B}}|}\biggl{\|}u^{k}-\sum_{j=1}^{r}f_{j}g_{j}\biggr{\|}_{L^{2}(X\times Y)}<\varepsilon,$ (4)
[…]
where, $\hat{\mathcal{B}}=\{\hat{\mathcal{B}}_{1},\hat{\mathcal{B}}_{2},\dots\hat{\mathcal{B}}_{k}\}$ is the set of sets consisting of active dimensions where at any given step-$k$, $\hat{\mathcal{B}}_{k}=\{\hat{\mathcal{B}}_{k}^{1},\hat{\mathcal{B}}_{k}^{2}\}=\{x_{p},x_{q}\}$ is the set of active dimensions, $\{X$$\hat{\mathcal{B}}_{k}^{1}\}$ and $\{Y$$\hat{\mathcal{B}}_{k}^{2}\}$ is the set of inactive dimensions. Here, $X=\{x_{1},x_{2},\dots x_{p}\dots x_{d/2}\}$, $Y=\{x_{d/2+1},\dots x_{q}\dots x_{d}\}$.
@@ 2.4 Effect of Preconditioning on Training Dynamics
where $f_{k}^{(\theta_{k})}(X_{k})=\sum_{i=1}^{n_{k}}w_{2,k}^{i}(w_{1,k}X_{k}+b_{1})=\sum_{i=1}^{n_{k}}w_{2,k}^{i}\phi_{i}(X_{k}),\leavevmode\nobreak\ k=1,2,3.$ The parameter $w^{i}_{j,k}$ denotes the weight associated with the $i^{\text{th}}$ neuron in the $j^{\text{th}}$ layer of the $k^{\text{th}}$ body network. In the above formulation, the basis functions are implicitly represented by the hidden layers of the body networks within the Anant-Net architecture, while the corresponding coefficients are extracted from the final output layer. This representation is analogous to the framework proposed by Cyr et al. [33], wherein a generic deep neural network is interpreted as a linear combination of learned basis functions. We consider the continuous formulation of the loss function, which serves as the analog of the discrete loss representation given in Equation (3), corresponding to the same PDE defined in Equation (2).
$\mathcal{J}(\theta)=\frac{1}{2}\int_{\Omega}|\mathcal{L}\hat{u}_{\theta}(x)-f(x)|^{2}dx+\frac{\lambda}{2}\int_{\partial\Omega}|\hat{u}_{\theta}(x)-g(x)|^{2}d\Omega,$ (6)
@@
where $H_{r}$ is the hessian computed at intermediate parameters between 0 and $r$. Here, for brevity a new notation $\phi_{i}(x)\triangleq\partial_{\theta_{i}}u(x;\theta_{0})$ is introduced with $\mathcal{L}\phi_{i}\in L^{2}(\Omega)$. The above formulation is substituted in Equation 6 that leads to the following simplification:
@@
$\mathbb{A}_{i,j}=\langle\mathcal{L}\phi_{i},\mathcal{L}\phi_{j}\rangle_{L^{2}{(\Omega)}}+\lambda\langle\phi_{i},\phi_{j}\rangle_{L^{2}{(\Omega)}}\hskip 28.45274pt\mathbb{A}\in\mathbb{R}^{n\times n},$
@@
In the case where the matrix $\mathbb{A}$ is invertible, let $\kappa(\mathbb{A})$ denote its condition number. To achieve a specified error tolerance $\epsilon$, the number of gradient descent iterations required satisfies the relation $N(\epsilon)\propto 1/\ln(1-c/\kappa(\mathbb{A}))$ for some constant $0<c<1$. Consequently, as $\mathbb{A}$ becomes increasingly ill-conditioned, i.e., $\kappa(\mathbb{A})\to\infty$, the number of iterations $N(\epsilon)$ diverges, indicating a deterioration in convergence rate. To address this challenge in the training of physics-informed machine learning models, we employ an explicit preconditioning strategy for the gradients, as detailed below.
@@
$\widehat{\theta}_{r+1}=\widehat{\theta}_{r}-\eta H^{-1}\nabla L(\widehat{\theta}_{r}).$ (9)
$\text{Hessian:}\hskip 14.22636ptH$ $=\partial_{\theta_{i}}\partial_{\theta_{j}}\mathcal{J}(\theta)$
@@ 2.5 Anant-Net with Interpretable KAN
$\Phi_{l}=\begin{bmatrix}\phi_{l,1,1}(\cdot)&\phi_{l,1,2}(\cdot)&\dots&\phi_{l,1,n_{l}}(\cdot)\\ \phi_{l,2,1}(\cdot)&\phi_{l,2,2}(\cdot)&\dots&\phi_{l,2,n_{l}}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{l,n_{l+1},1}(\cdot)&\phi_{l,n_{l+1},2}(\cdot)&\dots&\phi_{l,n_{l+1},n_{l}}(\cdot)\\ \end{bmatrix}$.
@@
Figure 4: A schematic of a 3-layer Kolmogorov–Arnold Network (KAN) is shown, with input $\mathbf{x}_{0}=\{x_{1},x_{2}\}$ and output $f(\mathbf{x})=(\Phi_{2}\circ\Phi_{1}\circ\Phi_{0})(\mathbf{x}_{0})$. Replacing the feedforward neural networks (body networks) in the Anant-Net architecture (as shown in Figure 2) with KAN-based body networks yields the Anant-KAN architecture.
@@ 3 Results
This section presents a series of results demonstrating the capability of Anant-Net in solving high-dimensional PDEs, both linear and nonlinear, as well as steady-state and transient, defined over hypercubic domains with side length $L>1$. Specifically, we focus on the Poisson equation, sine-Gordon equation, and Allen-Cahn equation. A key challenge in such settings is the enforcement of boundary conditions, where boundary sampling significantly contributes to the complexity of the problem as the dimensionality increases. To evaluate the performance of the proposed architecture, we use the percentage relative $L_{2}$ error, defined as:
@@
Uncertainty quantification is performed by reporting the test error over 10 independent random realizations, with data points sampled from the high-dimensional domain. Unless otherwise specified, all experiments are executed on a T4 GPU with 15 GB RAM. To this end, we systematically compare the performance of Anant-Net with other state-of-the-art methods that also employ boundary sampling strategies, rather than enforcing boundary conditions directly. The weights for the individual loss terms is fixed for all the cases: $\lambda_{r}=1$ and $\lambda_{b}=15$ in Equation 3. We use Xavier initialization to initialize the trainable parameters of the model for all the cases. For training the model, we use both AdamW and L-BFGS optimizers with learning rates, 1e031e-03 and 1e021e-02 respectively.
In addition to the architectural hyperparameters in deep neural networks, several algorithmic hyperparameters must be tuned to achieve optimal performance with the Anant-Net architecture. These include, the number of body networks ($B$), grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#CG\#CG), number of boundary/data grids (#BG\#BG), and the sampling frequency ($S_{\text{FREQ}}$). In a way, the total number of iterations for training the model is approximately equal to $S_{\text{FREQ}}\times\#CG$ and the training is stopped when the maximum number of iterations is reached. The number of body networks ($B$) directly influences the architecture of Anant-Net. For example, when $B=3$, the architecture consists of three body networks that process 3D cubes sampled from the high-dimensional space. Similarly, when $B=4$, four body networks are used to process 4D tesseracts from the high-dimensional domain. The sampling frequency ($S_{\text{FREQ}}$) directly affects the training process by determining how frequently a new set of dimensions is sampled during training. If $S_{\text{FREQ}}$ is too low, the model spends less time interacting with the newly sampled subspace, leading to poor generalization. Conversely, if $S_{\text{FREQ}}$ is too high, the model may overfit due to prolonged exposure to the same subspace. Finally, the grid resolution for collocation points ($N_{C}$), grid resolution for boundary/data points ($N_{B}$), number of collocation grids (#CG\#CG), and number of boundary/data grids (#BG\#BG) determine the size and structure of the dataset handled by Anant-Net. We will employe $B=3$ for all computational experiments unless stated otherwise.
Figure 5: Point-wise absolute testing error on randomly selected 3D cubes from 21D, 99D, and 300D cases for the (left) Poisson equation, (center) sine-Gordon equation, and (right) Allen–Cahn equation. Here, $\mathsf{ActiveDims}$: $\{p,q,r\}$” denotes the active dimensions in the $d$-dimensional space where gradients are non-zero, and $\mathrm{D}\{$$\{p,q,r\}$ represents the inactive dimensions. $\mathrm{D}=\{1,2,\dots,d\}$ denotes the set of all dimensions.
@@ 3.1 High-dimensional Poisson’s Equation
Dimension Architecture Optimizer #\#Collocation #\#Data
@@
In this section, we solve the high-dimensional Poisson’s equation in 21D, 99D, and 300D. For all cases, we use the Adam optimizer to train Anant-Net, followed by L-BFGS to further fine-tune the learning process. The benefits of preconditioning the gradients were discussed in previous sections, where we empirically validated the theory using (quasi-)Newton optimization with a linear version of Anant-Net. However, in this section, Anant-Net incorporates a non-linear activation function ($\tanh$) to solve the high-dimensional Poisson’s equation. Table 2 provides further details on the specific hyperparameters used. As suggested by the curse of dimensionality, the number of collocation and data samples required to solve high-dimensional PDEs on a hypercube grows exponentially with dimension. This poses a major challenge for existing methods, including state-of-the-art PINNs such as NTK-based PINNs [30], self-adaptive PINNs [40], etc., which struggle with scalability due to the computational burden of large sample volumes. Anant-Net addresses this issue efficiently. Table 2 summarizes the total number of collocation and boundary points used for solving Poisson’s equation in 21D, 99D, and 300D. The sample sizes are controlled via hyperparameters: $N_{C}$, #CG\#CG for collocation points, and $N_{B}$, #BG\#BG for boundary points. Figure 6 (left) indicates that boundary data sampling is the primary contributor to training time across all dimensions, highlighting a need to optimize this component. Meanwhile, Figure 6 (right) shows a mean relative $L_{2}$ testing error along with the standard deviation, evaluated on randomly selected test points from the high-dimensional space. Figure 5 (left) shows the absolute point-wise error on randomly selected 3D cubes from the 21D, 99D, and 300D cases. More details on the hyperparameters used for training is explained in APPENDIX B.